a car is traveling on a straight road at 19 m/s. the driver sees a bird standing in the road 38 meters ahead and applies the brakes, giving constant deceleration of 5 m/s2.

how far away is the driver from the bird when he stops( assuming the bird doesn't move)


1
Expert's answer
2021-11-24T11:50:10-0500

"v=0" m/s , when he stops

"v_0=19" m/s, initial velocity

"a=-5" m/s"^2" , constant deceleration

"\Delta x=\frac{v^2-v_0^2}{2a}=\frac{0^2-19^2}{2\cdot (-5)}=36.1" m


"38-36.1=1.9" m


Answer: "1.9" m from the bird.